7. A survey of 800 college students is conducted to see whether they are a new student this
quarter or not. The random variable respresents the number of students enrolled as a new student. Is
this a binomial experiment? Why or why not?

Yes or No (Circle One)


8. A recent survey found that 79% of all adults over 50 wear sunglasses for driving. In a random sample of 80 adults over 50, what is the mean and standard deviation of those that wear sunglasses?

Mean :__________

Standard Deviation :________
Answer #1

Question 7

Answer:

Yes, because this experiment has only two outcomes as new or old.

Question 8

We are given n = 80, p = 0.79, q = 1 – p = 1 – 0.79 = 0.21

Mean = np = 80*0.79 = 63.2

Mean = 63.2

Standard deviation = sqrt(npq) = sqrt(80*0.79*0.21) = 3.643076

Standard deviation = 3.643076